Hello, I need help with my trigonometry question from my study guide.
Answer #1

since here Ô is measured in degree, so we have 14 = CO * 20 * 30 >> A = => 0 = 14X3600 2 TX30 7 * 12° >>0 = 890 - O = 126.79